Theory of Interest: The sale price of the car you wanted to buy is $25,456, but you decided to buy the car for a deal of 60 monthly payments of $559 each month. After a year, you decide it's time for some new wheels. You return to the dealer and want to trade in your car for a new model. The dealer offers you a trade-in value of $20,000. How much of a down payment will you be able to make after you paid off the remainder of your loan on your first car? Please explain thoroughly and if you are using Excel worksheets, please write the formulas you used to solve this problem.
Answer #1

Answer:

Car price = PV = $25,456

Monthly Payment = PMT = $559

Number of monthly payments = NPER = 60

Monthly interest rate = RATE(nper, pmt, pv, fv, type) = RATE (60, 559, -25456, 0, 0) = 0.95282858255%

Let us calculate outstanding balance after 1 year (12 months)

Outstanding balance = Present value of remaining monthly installments

Remaining number of monthly installment after one year = 60 - 12 = 48

Outstanding balance after one year = PV(rate, nper, pmt, fv, type) = PV ( 0.95282858255%, 48, -559,0, 0) = $21,453.16

As such:

Trade-in value of the car = $20,000 and

Outstanding balance after one year = $21,453.16

So to settle the remainder of the loan you have to pay more than what you get from trade-in value of the car.

As such:

Down payment you be able to make after you paid off the remainder of your loan on your first car = $0